Suppose the total product (Q) is given by the production function; [Q=4KL^-0.5] where K=capital=12 and L= unit of labor = -36. Calculate the adverage product of labor.


1
Expert's answer
2021-03-14T18:56:44-0400
"\frac {\delta Q}{\delta L}=4 \times (-0.5) \times KL^{-1.5}"

If K=12, L=36, then


"\frac {\delta Q}{\delta L}=-2\times 12 \times (6^2)^{-1.5}=-\frac {1}{9}"